OpenStudy (anonymous):

The sum of two consecutive integers is between 12 and 16. What are the two integers? A. 13 and 14 B. 7 and 8 C. 14 and 15

Parth (parthkohli):

\(\Large \color{midnightblue}{\rightarrow 13 + 14 = 17 \rightarrow nope }\) \(\Large \color{midnightblue}{\rightarrow 7 + 8 = 15 \rightarrow \checkmark }\) \(\Large \color{midnightblue}{\rightarrow 14 + 15 = 29 \rightarrow nope }\)